OpenStudy (anonymous):

When Mr Johnson was over 50 but less than 80 he told a friend: "Each of my sons has as many sons as brothers, and the number of my years is exactly that of the number of both my sons and grandsons." How many sons and grandsons had Mr Johnson?

OpenStudy (anonymous):

is this harder than i think? or am i making a mistake?

OpenStudy (anonymous):

I think it's pretty hard :D

OpenStudy (anonymous):

i am thinking you can guess and check for this one

OpenStudy (anonymous):

say five sons then each son has 4 children so there are 5 sons and 20 grandsons for a total of 25 ok that doesn't work, try a bigger number

OpenStudy (anonymous):

Bingo! Eight sons and 56 grandsons. So he was 64, right?

OpenStudy (anonymous):

\[8\times 7+8=8^2=64\] yup

OpenStudy (anonymous):

if you want to sound fancy you can say that 64 is the only perfect square between 50 and 80 since \(7^2=49\) and \(9^2=81\)

OpenStudy (anonymous):

Thank you :)